# Cognitive Visual-Language Mapper: Advancing Multimodal Comprehension with Enhanced Visual Knowledge Alignment

### v1

@@ Abstract
Evaluating and Rethinking the current landscape of Large Multimodal Models (LMMs), we observe that widely-used visual-language projection approaches (e.g., Q-former or MLP) focus on the alignment of image-text descriptions yet ignore the visual knowledge-dimension alignment, i.e., connecting visuals to their relevant knowledge. Visual knowledge plays a significant role in analyzing, inferring, and interpreting information from visuals, helping improve the accuracy of answers to knowledge-based visual questions. In this paper, we mainly explore improving LMMs with visual-language knowledge alignment, especially aimed at challenging knowledge-based visual question answering (VQA). To this end, we present a Cognitive Visual-Language Mapper (CVLM), which contains a pretrained Visual Knowledge Aligner (VKA) and a Fine-grained Knowledge Adapter (FKA) used in the multimodal instruction tuning stage. Specifically, we design the VKA based on the interaction between a small language model and a visual encoder, training it on collected image-knowledge pairs to achieve visual knowledge acquisition and projection. FKA is employed to distill the fine-grained visual knowledge of an image and inject it into Large Language Models (LLMs). We conduct extensive experiments on knowledge-based VQA benchmarks and experimental results show that CVLM significantly improves the performance of LMMs on knowledge-based VQA (average gain by 5.0%). Ablation studies also verify the effectiveness of VKA and FKA, respectively.
@@ 1 Introduction
Recent Large Multimodal Models (LMMs) such as GPT-4V OpenAI (2023), Gemini Team et al. (2023), MiniGPT-4 Zhu et al. (2023), InstructBLIP Liu et al. (2023b), LLaVA Liu et al. (2023a), and many others, have achieved impressive performance in a variety of visual understanding and reasoning tasks, especially on Visual Question Answering (VQA) Li et al. (2023d, f). Current open-source LMMs are usually constructed by combining pertained visual encoders and Large Language Models (LLMs), as depicted in the bottom part of Figure 1, where a visual mapping network (e.g., Q-former Li et al. (2023c), Linear Zhu et al. (2023), or MLP Liu et al. (2023a); Li et al. (2023e)) is employed to project visual representations into the language space of LLMs. Although such LMMs have achieved powerful visual understanding capability similar to GPT-4V and Genimi on some image understanding tasks such as Image Captioning Changpinyo et al. (2021), Visual Dialogue Zhang et al. (2022); Chen et al. (2022), Visual Entailment Xie et al. (2019); Do et al. (2020), and VQA Antol et al. (2015), they often fall short of knowledge-based VQA, which necessitates relevant knowledge to answer these visual questions. As the cases illustrated in Figure 1, these advanced LMMs (including GPT-4V and Gemini-Pro) can not give correct answers to simple visual information seeking questions: Who is the manufacturer of this aircraft; What country does this building belong to?.
In light of this, rethinking the construction process of LMMs Wang et al. (2022); Li et al. (2023f); Zhu et al. (2023); Liu et al. (2023a); Koh et al. (2023) from the initial pretraining stages, we discover that these visual mapping networks trained on massive image-text captioning pairs simply transfer visual features to their language descriptions. They overlook the visual language knowledge-dimension alignment. i.e., connecting visuals to their relevant knowledge. As we know, visual knowledge Collins and Olson (2014) plays a pivotal role in the way humans understand and interact with the world. It extends beyond the mere ability to recognize and interpret visuals, incorporating an understanding of spatial relationships, patterns, and symbols, which are essential components of human cognition Pinker (1984); Cavanagh (2011). Additionally, previous works also demonstrated that introducing visual knowledge Lu et al. (2022); Zhu et al. (2022); Li et al. (2023e, g) can improve the performance of pretrained language models on natural language understanding (Lu et al., 2022) and open-ended text generation tasks (Zhu et al., 2022). Inspired by these insights, we focus on enhancing LMMs through the introduction of visual-language knowledge alignment, going beyond the conventional scope of visual-language integration.
@@
  • We present a cognitive visual-language mapper to achieve visual-language knowledge alignment, which contains a pretrained visual knowledge aligner and a fine-grained knowledge adapter that is used to distill and inject valuable visual knowledge into LLMs.
@@ 2 Related Work
Large Visual-Language Models. Recent advancements in foundational models for vision and language have led to the development of LMMs. In response to large model GPT-4 OpenAI (2023), several others have emerged, including GVT Wang et al. (2023), MPlug Ye et al. (2023), Macaw Lyu et al. (2023), LMEye Li et al. (2023f), and LLaVA Liu et al. (2023a). These models have demonstrated strong performance across various visual-language tasks. Typically, these models utilize pretrained visual models to extract visual features, which are then integrated into the linguistic space of LMMs through a straightforward projection layer. This layer can be a Linear Layer Merullo et al. (2023); Liu et al. (2023b); Li et al. (2023e) or a Q-former Li et al. (2023c). Following this integration, similar to the supervised fine-tuning approach used for LLMs, these systems undergo refinement using diverse and high-quality multimodal instruction-following datasets Liu et al. (2023b); Zhu et al. (2023); Ye et al. (2023). These datasets encompass both human-labeled data for downstream tasks like Visual Question Answering (VQA) and VCR Antol et al. (2015); Zellers et al. (2019), as well as datasets automatically generated by GPT-4. Meanwhile, some multimodal benchmarks like MMBench Liu et al. (2023c) and SEED-Bench Li et al. (2023a) have been established to evaluate advanced LMMs, and Li et al. (2023h) presents a comprehensive assessment of their performance in knowledge-intensive VQA scenarios.
@@ 3.2 Visual Knowledge Aligner
$\begin{array}[]{c}\mathbf{h}_{\text{AO}}=\text{VKA}([\mathbf{h}_{\text{KP}},\mathbf{h}_{\text{KQ}}],\mathbf{h}_{I}),\\ \mathbf{h}_{\text{KO}}=\mathbf{W}^{K}\mathbf{h}_{\text{AO}}+\mathbf{b},\end{array}$ (1)
@@ 3.3 Fine-grained Visual Knowledge Adapter
$\begin{array}[]{c}\mathbf{h}_{S}^{l-1}=\text{Self-A}(\mathbf{h}_{D},\mathbf{h}_{T})+\mathbf{h}_{\text{FKA}}^{l-1},\\ \mathbf{h}_{cr}=\text{Cross-A}(\mathbf{h}_{S}^{l-1},[\mathbf{h}_{I_{1}},\mathbf{h}_{I_{2}},...,\mathbf{h}_{I_{5}}])+\mathbf{h}_{S}^{l-1},\\ \mathbf{h}_{\text{FKA}}^{l}=\text{MLP}(\mathbf{h}_{cr}),\end{array}$ (2)
@@ 3.4 Training
$\begin{array}[]{c}\mathcal{L}_{\text{VKA}}^{P}=-\sum_{i=1}^{\text{KM}}\mathbf{log}{P}_{i}(\hat{y}_{i}=y_{i}|h_{I},h_{\text{KP}};\\ y_{1},...,y_{i-1}),\\ \mathcal{L}_{\text{VKA}}^{A}=-\sum_{i=1}^{\text{KM}}\mathbf{log}{P}_{i}(\hat{y}_{i}=y_{i}|h_{I},h_{\text{KP}},h_{\text{KQ}};\\ y_{1},...,y_{i-1}),\end{array}$ (3)
@@ 4 Experiments
${}^{\ddagger}$ Vicuna-7B 53.00 50.8 72.5 73.45 65.27 45.81 8.18 55.05
@@
CVLM (3M IKPairs) Vicuna-7B 57.83 56.92 76.48 78.95 70.39 50.48 11.27 60.30
Table 1: Comparison between different LMMs on knowledge-based VQA benchmarks. With 7B parameters, CVLM achieves the best performance with the same training data. $\ddagger$” shows that we fairly use the same instruction tuning data to train the model. “IKPairs” represents the image-knowledge pairs used to train VKA and the initial version is trained with 2M pairs. Benchmark names are abbreviated due to space limits. A-OKVQAM: Multi-Choice A-OKVQA Schwenk et al. (2022); A-OKVQAO: Open-ended A-OKVQA Schwenk et al. (2022); TextVQA Singh et al. (2019); Infoseek Chen et al. (2023b); SEED-BenchS: SEED-Bench (Spatial) Li et al. (2023b);
@@ 4.1 Data sets
${}^{\clubsuit}$ (Vicuna-13b) 8.50 3.3 2.0 1.33 10.0 10.67 6.0 4.67 26.67 2.67 20.67 5.33
${}^{\clubsuit}$ (FlanT5-xxl) 8.37 4.0 5.33 2.0 8.67 8.0 8.0 8.0 28.0 5.34 8.67 6.0
${}^{\clubsuit}$ 10.22 11.33 16.67 0.0 24.67 6.0 0.7 10.67 26.0 5.3 0.13 10.0
${}^{\ddagger\clubsuit}$ 8.18 5.33 6.67 3.33 10.00 11.33 6.67 3.33 28.67 2.67 5.33 6.67
@@
${}^{\clubsuit}$ (Vicuna-13b) 41.02 34.00 52.41 37.33 51.33 33.33 46.34 31.33 38.67 32.30 49.11 43.33
${}^{\clubsuit}$ (FlanT5-xxl) 47.96 44.66 51.03 48.67 48.0 43.33 51.22 47.33 42.0 55.38 58.04 45.33
${}^{\clubsuit}$ 57.25 50.0 62.76 58.0 62.67 54.0 60.16 50.0 53.33 61.54 65.18 57.33
${}^{\ddagger\clubsuit}$ 52.64 48.0 53.10 46.67 58.67 52.67 57.72 45.33 49.33 55.38 59.82 56.67
@@
Table 3: Held-In testing results on OK-VQA with Commonsense Knowledge. Baseline results are reported by Li et al. (2023h). Knowledge names are abbreviated due to space limits. and Transportation (VT); Brands, Companies and Products (BCP); Objects, Material and Clothing (OMC); Sports and Recreation (SR); Cooking and Food (CF); Geography, History, Language and Culture (GHLC); People and Everyday Life (PEL); Plants and Animals (PA); Science and Technology (ST); Weather and Climate (WC); and Other.
@@ 4.3 Implementation Details
We utilize the AdamW Kingma and Ba (2014) optimizer with a cosine learning rate scheduler to train our model. During the pretraining stage of VKA, we first train it on 2 A100 GPUs using a dataset of 2 million image-knowledge pairs from Wikipedia with a global batch size of 128 and a base learning rate of 5e-5. In the alignment stage, the model is trained on the same 2 million Wikipedia data using 2 A100 GPUs with a global batch size of 32 and a maximum learning rate of 1e-4. For the final stage, we employ LoRA to efficiently fine-tune the language model. In our implementation, we set the rank to 128 and alpha to 256, with a learning rate of 1e-4 for LoRA parameters and the newly added FKA. We use a smaller learning rate of 2e-5 for MLP and VKA.
@@ 4.4 Main Results
Overall Performance. We present the comparative performance of all models across seven benchmarks tailored for knowledge-based VQA tasks, as detailed in Table 1. Our method exhibits a significant improvement over established baselines, with performance gains of 4% and 4.5% on dataset A-OKVQA and TextVQA, respectively. These advancements underscore the effectiveness of our visual-knowledge alignment modules VKA and FKA in bolstering the capabilities of LMMs, particularly evident in the enhancements to ${}^{\ddagger}$. Despite these improvements, our model slightly underperforms in comparison to InstructBLIP on the SEED-Bench, which may be attributed to the larger scale of multimodal instruction tuning data and larger language models (FlanT5-XXL-11B) used by InstructBLIP.
@@ 4.5 Ablation Study
Effects of Visual Knowledge Aligner. To assess the impact of the Visual Knowledge Aligner on model performance, we trained LLaVA-v1.5 using the identical 504K dataset mentioned earlier. As depicted in Table 1, when compared to the baseline ${}^{\ddagger}$, CVLM(len=0) incorporating the Visual Knowledge Aligner yielded improved results across all benchmarks. Specifically, CVLM(len=0) exhibited the most significant enhancement on SEED-Bench(Spatial), achieving a 4.5% increase.
@@
Impact of the Size of Visual Knowledge Pairs. To validate the effectiveness of adding more knowledge pre-training data, we increased the amount of Wikipedia knowledge pre-training data from 2M to 3.3M during training VKA with LLM. The experimental results of CVLM (3M IKPairs) are shown in Tables 1, 2 and 3. We observed that the inclusion of more pretraining knowledge data significantly enhances the model’s ability to comprehend knowledge, consequently resulting in higher performance on various tasks.
Analysis of Distillation Vector Length. Our examination of the optimal distillation vector length, as shown in Tables 2 and 3, indicates that increasing distillation vector length does not significantly improve model performance, rather results in performance fluctuations. This suggests that expanding distillation vectors could disturb the structural integrity of large language models, potentially reducing our model’s effectiveness, especially in knowledge-dependent tasks.
Impact of introducing more instruction datasets. As the bottom results shown in Tables 2 and 3, we can see that more instruction data (624k from LLaVA-v1.5) will bring improvement on two knowledge-based VQA datasets, yet it will degrade the performance on some knowledge categories such as Plant. It indicates that introducing more data may not necessarily bring about an overall improvement in performance. Our cognitive mapper method leads to greater performance improvements with a small amount of instruction data, compared to adding more instruction data.

### v2

@@ Abstract
Evaluating and Rethinking the current landscape of Large Multimodal Models (LMMs), we observe that widely-used visual-language projection approaches (e.g., Q-former or MLP) focus on the alignment of image-text descriptions yet ignore the visual knowledge-dimension alignment, i.e., connecting visuals to their relevant knowledge. Visual knowledge plays a significant role in analyzing, inferring, and interpreting information from visuals, helping improve the accuracy of answers to knowledge-based visual questions. In this paper, we mainly explore improving LMMs with visual-language knowledge alignment, especially aimed at challenging knowledge-based visual question answering (VQA). To this end, we present a Cognitive Visual-Language Mapper (CVLM), which contains a pretrained Visual Knowledge Aligner (VKA) and a Fine-grained Knowledge Adapter (FKA) used in the multimodal instruction tuning stage. Specifically, we design the VKA based on the interaction between a small language model and a visual encoder, training it on collected image-knowledge pairs to achieve visual knowledge acquisition and projection. FKA is employed to distill the fine-grained visual knowledge of an image and inject it into Large Language Models (LLMs). We conduct extensive experiments on knowledge-based VQA benchmarks and experimental results show that CVLM significantly improves the performance of LMMs on knowledge-based VQA (average gain by 5.0%). Ablation studies also verify the effectiveness of VKA and FKA, respectively. [^1] Codes and Data are available at https://github.com/HITsz-TMG/Cognitive-Visual-Language-Mapper
@@ 1 Introduction
Recent Large Multimodal Models (LMMs) such as GPT-4V OpenAI (2023), Gemini Team et al. (2023), MiniGPT-4 Zhu et al. (2023), InstructBLIP Liu et al. (2023b), LLaVA Liu et al. (2023a), and many others, have achieved impressive performance in a variety of visual understanding and reasoning tasks, especially on Visual Question Answering (VQA) Li et al. (2023e, g). Current open-source LMMs are usually constructed by combining pertained visual encoders and Large Language Models (LLMs), as depicted in the bottom part of Figure 1, where a visual mapping network (e.g., Q-former Li et al. (2023c), Linear Zhu et al. (2023), or MLP Liu et al. (2023a); Li et al. (2023f)) is employed to project visual representations into the language space of LLMs. Although such LMMs have achieved powerful visual understanding capability similar to GPT-4V and Genimi on some image understanding tasks such as Image Captioning Changpinyo et al. (2021), Visual Dialogue Zhang et al. (2022); Chen et al. (2022), Visual Entailment Xie et al. (2019); Do et al. (2020), and VQA Antol et al. (2015), they often fall short of knowledge-based VQA, which necessitates relevant knowledge to answer these visual questions. As the cases illustrated in Figure 1, these advanced LMMs (including GPT-4V and Gemini-Pro) can not give correct answers to simple visual information seeking questions: Who is the manufacturer of this aircraft; What country does this building belong to?.
In light of this, rethinking the construction process of LMMs Wang et al. (2022); Li et al. (2023g); Zhu et al. (2023); Liu et al. (2023a); Koh et al. (2023) from the initial pretraining stages, we discover that these visual mapping networks trained on massive image-text captioning pairs simply transfer visual features to their language descriptions. They overlook the visual language knowledge-dimension alignment. i.e., connecting visuals to their relevant knowledge. As we know, visual knowledge Collins and Olson (2014) plays a pivotal role in the way humans understand and interact with the world. It extends beyond the mere ability to recognize and interpret visuals, incorporating an understanding of spatial relationships, patterns, and symbols, which are essential components of human cognition Pinker (1984); Cavanagh (2011). Additionally, previous works also demonstrated that introducing visual knowledge Lu et al. (2022b); Zhu et al. (2022); Li et al. (2023f, h) can improve the performance of pretrained language models on natural language understanding (Lu et al., 2022b) and open-ended text generation tasks (Zhu et al., 2022). Inspired by these insights, we focus on enhancing LMMs through the introduction of visual-language knowledge alignment, going beyond the conventional scope of visual-language integration.
@@
  • We present a cognitive visual-language mapper to achieve visual-language knowledge alignment, which contains a pretrained visual knowledge aligner and a fine-grained knowledge adapter that is used to distil and inject valuable visual knowledge into LLMs.
@@ 2 Related Work
Large Visual-Language Models. Recent advancements in foundational models for vision and language have led to the development of LMMs. In response to large model GPT-4 OpenAI (2023), several others have emerged, including GVT Wang et al. (2023), MPlug Ye et al. (2023), Macaw Lyu et al. (2023), LMEye Li et al. (2023g), and LLaVA Liu et al. (2023a). These models have demonstrated strong performance across various visual-language tasks. Typically, these models utilize pretrained visual models to extract visual features, which are then integrated into the linguistic space of LMMs through a straightforward projection layer. This layer can be a Linear Layer Merullo et al. (2023); Liu et al. (2023b); Li et al. (2023f) or a Q-former Li et al. (2023c). Following this integration, similar to the supervised fine-tuning approach used for LLMs, these systems undergo refinement using diverse and high-quality multimodal instruction-following datasets Liu et al. (2023b); Zhu et al. (2023); Ye et al. (2023). These datasets encompass both human-labeled data for downstream tasks like Visual Question Answering (VQA) and VCR Antol et al. (2015); Zellers et al. (2019), as well as datasets automatically generated by GPT-4. Meanwhile, some multimodal benchmarks like MMBench Liu et al. (2023c) and SEED-Bench Li et al. (2023a) have been established to evaluate advanced LMMs, and Li et al. (2023i) presents a comprehensive assessment of their performance in knowledge-intensive VQA scenarios.
@@ 3.2 Visual Knowledge Aligner
$\begin{array}[]{c}\mathbf{h}_{\text{AO}}=\text{VKA}([\mathbf{h}_{\text{KP}},\mathbf{h}_{\text{KQ}}],\mathbf{h}_{I}),\vspace{1.0ex}\\ \mathbf{h}_{\text{KO}}=\mathbf{W}^{K}\mathbf{h}_{\text{AO}}+\mathbf{b},\end{array}$ (1)
@@ 3.3 Fine-grained Visual Knowledge Adapter
$\begin{array}[]{c}\mathbf{h}_{S}^{l-1}=\text{Self-A}(\mathbf{h}_{D},\mathbf{h}_{T})+\mathbf{h}_{\text{FKA}}^{l-1},\vspace{1.0ex}\\ \mathbf{h}_{cr}=\text{Cross-A}(\mathbf{h}_{S}^{l-1},[\mathbf{h}_{I_{1}},\mathbf{h}_{I_{2}},...,\mathbf{h}_{I_{5}}])+\mathbf{h}_{S}^{l-1},\vspace{1.0ex}\\ \mathbf{h}_{\text{FKA}}^{l}=\text{MLP}(\mathbf{h}_{cr}),\end{array}$ (2)
@@ 3.4 Training
$\begin{array}[]{c}\mathcal{L}_{\text{VKA}}^{P}=-\sum_{i=1}^{\text{KM}}\mathbf{log}{P}_{i}(\hat{y}_{i}=y_{i}|h_{I},h_{\text{KP}};\\ y_{1},...,y_{i-1}),\vspace{1.0ex}\\ \mathcal{L}_{\text{VKA}}^{A}=-\sum_{i=1}^{\text{KM}}\mathbf{log}{P}_{i}(\hat{y}_{i}=y_{i}|h_{I},h_{\text{KP}},h_{\text{KQ}};\\ y_{1},...,y_{i-1}),\end{array}$ (3)
@@ 4 Experiments
Qwen-VL Bai et al. (2023) Qwen 55.92 57.13 77.89 70.04 67.25 41.94 15.21 62.41
LLaVA-v1.5 Vicuna-7B 53.00 50.8 72.5 73.45 65.27 45.81 8.18 55.05
@@
CVLM (3M IKPairs, Objects=1) Vicuna-7B 57.92 56.90 76.32 78.69 70.48 50.21 10.01 59.65
CVLM (3M IKPairs, Objects=3) Vicuna-7B 58.19 57.17 76.40 79.21 70.91 50.32 10.42 62.93
CVLM (3M IKPairs Objects=5) Vicuna-7B 57.83 56.92 76.48 78.95 70.39 50.48 11.27 60.30
CVLM (3M IKPairs, Objects=8) Vicuna-7B 57.28 56.71 76.0 78.95 69.08 49.59 10.78 59.83
CVLM Qwen-VL 60.23 58.91 80.88 82.71 72.14 45.02 15.45 66.28
Table 1: Comparison between different LMMs on knowledge-based VQA benchmarks. With 7B parameters, CVLM achieves the best performance with the same training data. $\ddagger$” shows that we fairly use the same instruction tuning data to train the model. “IKPairs” represents the image-knowledge pairs used to train VKA and the initial version is trained with 2M pairs. “Objects” refers to the number of object regions used in FKA, which are obtained by SAM. Benchmark names are abbreviated due to space limits. A-OKVQAM: Multi-Choice A-OKVQA Schwenk et al. (2022); A-OKVQAO: Open-ended A-OKVQA Schwenk et al. (2022); TextVQA Singh et al. (2019); Infoseek Chen et al. (2023b); SEED-BenchS: SEED-Bench (Spatial) Li et al. (2023b);
@@ 4.1 Data sets
InstructBLIP (Vicuna-13b) 8.50 3.3 2.0 1.33 10.0 10.67 6.0 4.67 26.67 2.67 20.67 5.33
InstructBLIP (FlanT5-xxl) 8.37 4.0 5.33 2.0 8.67 8.0 8.0 8.0 28.0 5.34 8.67 6.0
LLaVA-v1.5-13b 10.22 11.33 16.67 0.0 24.67 6.0 0.7 10.67 26.0 5.3 0.13 10.0
LLaVA-v1.5-7b‡♣ 8.18 5.33 6.67 3.33 10.00 11.33 6.67 3.33 28.67 2.67 5.33 6.67
@@
Model MMbench ScienceQA-I PoPE-Adversarial PoPE-Random PoPE-Popular
CVLM w/o (FKA & VKA) 50.85 62.76 82.94 86.98 86.14
CVLM w/o FKA 56.46 69.21 83.13 87.81 85.85
CVLM 59.78 68.92 82.78 88.79 85.64
CVLM (3M IKPairs) w/o FKA 57.06 69.96 82.32 85.75 84.26
CVLM (3M IKPairs) 59.10 69.85 82.47 86.90 84.68
Table 3: Comparison of different models across various common benchmarks. ScienceQA Lu et al. (2022a) and PoPE Li et al. (2023d) are used to evaluate the complex reasoning and hallucination recognition abilities of LMMs.
@@
InstructBLIP (Vicuna-13b) 41.02 34.00 52.41 37.33 51.33 33.33 46.34 31.33 38.67 32.30 49.11 43.33
InstructBLIP (FlanT5-xxl) 47.96 44.66 51.03 48.67 48.0 43.33 51.22 47.33 42.0 55.38 58.04 45.33
LLaVA-v1.5-7b 57.25 50.0 62.76 58.0 62.67 54.0 60.16 50.0 53.33 61.54 65.18 57.33
LLaVA-v1.5‡♣ 52.64 48.0 53.10 46.67 58.67 52.67 57.72 45.33 49.33 55.38 59.82 56.67
@@
Table 4: Held-In testing results on OK-VQA with Commonsense Knowledge. Baseline results are reported by Li et al. (2023i). Knowledge names are abbreviated due to space limits. and Transportation (VT); Brands, Companies and Products (BCP); Objects, Material and Clothing (OMC); Sports and Recreation (SR); Cooking and Food (CF); Geography, History, Language and Culture (GHLC); People and Everyday Life (PEL); Plants and Animals (PA); Science and Technology (ST); Weather and Climate (WC); and Other.
@@ 4.3 Implementation Details
We utilize the AdamW Kingma and Ba (2014) optimizer with a cosine learning rate scheduler to train our model. During the pretraining stage of VKA, we first train it on 2 A100 GPUs using a dataset of 2 million image-knowledge pairs from Wikipedia with a global batch size of 128 and a base learning rate of 5e-5. Image knowledge processing is shown in the Appendix. In the alignment stage, the model is trained on the same 2 million Wikipedia data using 2 A100 GPUs with a global batch size of 32 and a maximum learning rate of 1e-4. For the final stage, we employ LoRA to fine-tune the language model efficiently. In our implementation, we set the rank to 128 and alpha to 256, with a learning rate of 1e-4 for LoRA parameters and the newly added FKA. We use a smaller learning rate of 2e-5 for MLP and VKA.
@@ 4.4 Main Results
Overall Performance. We present the comparative performance of all models across seven benchmarks tailored for knowledge-based VQA tasks, as detailed in Table 1. Our method exhibits a significant improvement over established baselines, with performance gains of 4% and 4.5% on dataset A-OKVQA and TextVQA, respectively. These advancements underscore the effectiveness of our visual-knowledge alignment modules VKA and FKA in bolstering the capabilities of LMMs, particularly evident in the enhancements to LLaVA-v1.5. Despite these improvements, our model slightly underperforms in comparison to InstructBLIP on the SEED-Bench, which may be attributed to the larger scale of multimodal instruction tuning data and larger language models (FlanT5-XXL-11B) used by InstructBLIP. We also evaluated our model on PoPE and ScienceQA in Table 3. On the ScienceQA dataset, we observe that including VKA results will largely improve the model performance yet FKA may not bring improvement. This may be attributed to the fine-grained features of images in ScienceQA containing less useful information. PoPE is an object recognition benchmark (only need the answer yes or no, like ”Is there an apple in the image”) and the similar performance indicates that introducing VKA and FKA will not affect the basic perception of an image.
@@ 4.5 Ablation Study
Effects of Visual Knowledge Aligner. To assess the impact of the Visual Knowledge Aligner on model performance, we trained LLaVA-v1.5 using the identical 504K dataset mentioned earlier. As depicted in Table 1, when compared to the baseline LLaVA-v1.5, CVLM(len=0) incorporating the Visual Knowledge Aligner yielded improved results across all benchmarks. Specifically, CVLM(len=0) exhibited the most significant enhancement on SEED-Bench(Spatial), achieving a 4.5% increase.
@@
Impact of the Size of Visual Knowledge Pairs. To validate the effectiveness of adding more knowledge pre-training data, we increased the amount of Wikipedia knowledge pre-training data from 2M to 3.3M during training VKA with LLM. The experimental results of CVLM (3M IKPairs) are shown in Tables 1, 2 and 4. We observed that the inclusion of more pretraining knowledge data significantly enhances the model’s ability to comprehend knowledge, consequently resulting in higher performance on various tasks.
Analysis of Distillation Vector Length. Our examination of the optimal distillation vector length, as shown in Tables 2 and 4, indicates that increasing distillation vector length does not significantly improve model performance, rather results in performance fluctuations. This suggests that expanding distillation vectors could disturb the structural integrity of large language models, potentially reducing our model’s effectiveness, especially in knowledge-dependent tasks.
Impact of introducing more instruction datasets. As the bottom results shown in Tables 2 and 4, we can see that more instruction data (624k from LLaVA-v1.5) will bring improvement on two knowledge-based VQA datasets, yet it will degrade the performance on some knowledge categories such as Plant. It indicates that introducing more data may not necessarily bring about an overall improvement in performance. Our cognitive mapper method leads to greater performance improvements with a small amount of instruction data, compared to adding more instruction data.
Analysis of the additional computation costs from SAM. We experimented with 1, 3, 5, and 8 image objects to determine the optimal count that maintains model performance while minimizing computational demand, as shown in Table 1. Our findings suggest that extracting 3 or 5 regions offers a balanced approach. Increasing the number of image regions tends to introduce visual noise, which can either slightly degrade performance or offer negligible improvements. Furthermore, most evaluation samples require limited fine-grained visual information, featuring fewer objects.